Lilly2010:

An expression to convert 50 miles per hour to miles per minute is shown. 50miles/1 hour x 1 hour/ ? minutes What value can be entered in the box to correctly make this conversion? 6 60 360 3,600

kristen000021:

60....

Luigi0210:

Well, how many minutes are in one hour?

kristen000021:

lilly is a bit slowww

Lilly2010:

@kristen000021 wrote:
lilly is a bit slowww
I am ya'll... and hanks lmao

kristen000021:

'hanks'

surjithayer:

1hour=60 minutes in 60 minutes distance covered=50 distance covered in 1 minute=50/60=5/6 miles

GRIZZLEY19:

The value that should be entered in the box is 60.